# Observational Probes of the Neutron Star Equation of State with Hyperons, Bosonic Dark Matter, and Quark Matter

### v2

@@ 3 EOS Characterization Across Phases and Thermodynamic Behavior
Figure 1: Pressure as a function of baryonic chemical potential for both hadronic phase and quark phase. The solid lines correspond to the hadronic phase with S for which the coupling strength of S remains constant ($x_{\text{S}}=0.03$) while different values of the mass of S are supposed. The QM EOSs shown with dotted lines are characterized by three key parameters: the diquark coupling ($\eta_{D}$), the vector meson coupling ($\eta_{V}$), and the square of the speed of sound, respectively. The hadronic EOS without S is also shown with the blue dashed line as DD2Y-T for comparison. The colored dots show the critical chemical potential in Eq. (8) at which the RIC has been applied.
@@
This study aims to explore the upper limit for the mass of S as a candidate for DM within the framework of the DD2Y-T model. To determine the allowed range for the mass of S, it is important to explore the behavior of matter at high densities and make sure the results agree with the observational constraints from NSs. It is essential to note that a mass shift of S particles, $x_{\text{S}}$, serves the purpose of mimicking the interaction with the surrounding medium. Given the postulation that S is a candidate for DM, it is reasonable to anticipate a predominantly weak interaction with the medium. Consequently, even though increasing the mass of S would inherently enable us to increase the value of $x_{\text{S}}$ based on our previous study (Shahrbaf et al., 2022), it is a deliberate choice to maintain $x_{\text{S}}$ as the coupling strength of the DM interaction with BM at its minimal value, specifically $x_{\text{S}}=0.03$ throughout the subsequent analyses. Notably, Moore & Slatyer (2024) highlights that for S particles to be a major DM candidate, their interactions must be extremely suppressed. This approach enables a systematic variation of the mass of the DM candidate, allowing us to examine its impact on the consistency of the EOS for hybrid stars, including hyperons in the hadronic phase, with current observational constraints.
@@
Therefore, we apply RIC, as outlined in Sect. 2.3, to model the phase transition from hadronic matter to deconfined QM. Within this approach, we study the influence of the mass of S particles on the hybrid EOS and aim to determine the maximum allowed value for the mass of S.
@@
Fig. 1 illustrates the pressure as a function of the baryonic chemical potential, showcasing both the hadronic phase and the quark phase. In the context of the hadronic phase, the coupling strength of S remains constant, as previously discussed, while different values of the mass of S with an increasing trend are supposed. One should notice that in comparison to the hadronic EOS without S (as represented by DD2Y-T, which includes nucleons and hyperons), the inclusion of S as a DM candidate results in an obvious softening of the EOS, although a repulsive interaction between DM and BM is considered. In the quark phase, we depict two various EOSs with specific parameters determined through BA, as outlined in reference Shahrbaf et al. (2023). Among the QM EOSs with the highest posterior probabilities in our BA, we specifically selected those that yield the earliest phase transition when combined with each hadronic EOS. These QM EOSs are characterized by three key parameters: the diquark coupling ($\eta_{D}$), the vector meson coupling ($\eta_{V}$), and the square of the speed of sound ($c_{s}^{2}=\frac{dP}{d\varepsilon}$), which serves as an indicator of EOS stiffness.
@@
The color dots show the crossing point between hadronic and quark lines, where the RIC with negative value of $\Delta_{P}$ in Eq. (9) has been applied to correct the mischaracterization of a transition from hadronic matter to deconfined QM. As a result, at lower densities, the hadronic phase dominates, while at high densities, the quark phase remains a valid description. One should also point out that the RIC construction with $\Delta_{P}<0$ generates a peak in the squared speed of sound, consistent with the phenomenologically established behavior observed in QCD-based studies.
@@
It is important to emphasize again that, for each hadronic EOS, we have selected the first intersection point with the QM EOS as the basis for implementing the phase transition construction. Although the two QM EOSs used in this study differ only slightly in stiffness, the softest three hadronic EOSs, corresponding to $m_{\text{S}}=$ 1890, 1895, and 1900 MeV, do not intersect with the stiffer of the two quark EOSs. Therefore, in the next section, where we perform a detailed scan over the mass of the S particle, these three soft hadronic models are combined with the first QM EOS (characterized by $\eta_{D}=0.70$, $\eta_{V}=0.10$, and $C_{s}^{2}=0.44$). The remaining hadronic models are matched with the second, slightly stiffer QM EOS (with $\eta_{D}=0.70$, $\eta_{V}=0.11$, and $C_{s}^{2}=0.44$).
@@
In Fig. 2, we illustrate how a RIC construction is performed between the hadronic EOS - represented by DD2Y-T+$\text{S}_{1900}$, which includes nucleons, hyperons, and a S with $m_{\text{S}}=1900$ MeV - and the QM EOS described by the nl-NJL model with parameters $\eta_{D}=0.70$, $\eta_{V}=0.10$, and $C_{s}^{2}=0.44$. The crossover phase transition between the two phases is characterized by a negative value of $\Delta_{P}$. The matching points defined earlier as $\mu_{H}$ and $\mu_{Q}$, representing the chemical potentials at which the pressure of the crossover phase is in concordance with the pressures of the hadronic and QM phases, is clearly illustrated in the figure. This figure shows how a negative value for $\Delta_{P}$ in the context of RIC effectively addresses the issue of a nonphysical intersection point, often referred to as the reconfinement point. This adjustment restores the hadronic phase at lower densities and the QM phase at higher densities.
@@
Figure 3: Fraction of S particles as a DM candidate in the hadronic phase, calculated using the DD2Y-T model for various S particle masses. The fractions are shown from the respective onset densities up to densities beyond the central density of NSs, illustrating how the S fraction evolves within the purely hadronic phase. The black and red dashed lines indicate the onset densities for the transition to deconfined QM for $m_{S}=1890$ MeV and $m_{S}=2054$ MeV, respectively.
@@
Regarding the production of S particles inside the hybrid star modeled by the approach described above, Fig. 3 shows the variation of the DM particle fraction as a function of baryon number density for different S-particle masses. In general, lighter S particles result in a higher DM fraction at a given density. Increasing the mass of S shifts the DM onset to higher densities, leading to a smaller fraction of heavier S particles being produced in stellar matter. For a fixed S mass, the DM fraction increases as the system becomes denser. However, at high densities - around $n_{B}\approx 1fm^{-3}$ the DM fraction becomes roughly the same across all considered S masses shown in the legend. The vertical black and red dashed lines indicate the deconfinement onset densities for $m_{S}=1890$ MeV and $m_{S}=2054$ MeV, respectively, marking the phase transition to QM. Notably, the maximum DM fraction for each mass is reached at this threshold density. Although the DM fractions differ across the curves at intermediate densities, the maximum S fraction inside the NS remains approximately $16\%$ for all masses. This behavior reflects that increasing the DM mass not only shifts the DM onset to higher densities but also delays the deconfinement transition, pushing it to higher densities as well, as it was also seen in Fig. 1.
4 Neutron Star Constraints on the Mass of DM Particles
@@
Figure 4: Mass-radius (M-R) relations for NSs with a purely hadronic core (blue dashed line), hybrid stars without DM (red dash-dotted line), and hybrid stars including DM for different S particle masses (solid colored lines). The colored regions represent observational constraints for comparison. The 1-$\sigma$ M-R constraint from the NICER analysis of PSR J0740+6620 is shown in turquoise (Dittmann et al., 2024a). The updated $95\%$ and $68\%$ credible NICER results for PSR J0030+0451 are displayed as nested light and dark cyan regions (Vinciguerra et al., 2023). The hatched magenta and green regions show the high-mass BW pulsar PSR J0952-0607 (Romani et al., 2022) and PSR J0348+0432 (Antoniadis et al., 2013), respectively. The credible intervals for HESS J1731-347 are shown as nested green regions, representing the $68\%$ (inner) and $90\%$ (outer) levels based on X-ray spectral modeling (Doroshenko et al., 2023). The M-R constraints for PSR J0437-4715 are represented by overlapping dark and light blue regions, corresponding to the $68\%$ (inner) and $90\%$ (outer) credible intervals (Choudhury et al., 2024a). The nested light and dark maroon regions show the most recent NICER analysis for the mass and radius of PSR J0614-3329, $95\%$ and $68\%$ credible results, respectively (Mauviard et al., 2025a). The zoomed-in region (a) highlights that all RIC EOSs including S DM pass through the $95\%$ credible region of PSR J0614–3329, whereas the RIC_DD2Y-T and DD2Y-T curves do not satisfy this constraint. The zoomed-in region (b) illustrates that for S masses larger than $1930$ MeV, the NS radius lies outside the $68\%$ credible region for PSR J0437–4715.
@@
As explained in the previous sections, recent multimessenger observations of NSs offer a valuable chance to explore the structure of high-density matter, which may contain exotic components such as hyperons, QM, or even DM. In this section, based on the astrophysical constraints of NSs, we investigate our hybrid EOS and determine the allowed maximum mass for the S particle. For this purpose, the latest M-R measurements by NICER telescope have been applied, which provide reliable constraints for the radius of NSs with masses around $2M_{\odot}$ and $1.4M_{\odot}$ (Miller et al., 2019; Riley et al., 2019; Miller et al., 2021; Riley et al., 2021; Choudhury et al., 2024a; Mauviard et al., 2025b). Since all of our hybrid EOSs pass well through the center of the NICER constraint region for PSR J0740+6620 with a mass of $2M_{\odot}$, we now focus primarily on the most recent results for the pulsars PSR J0614-3329 which has a measured mass of $M=1.44^{+0.06}_{-0.07}\,M_{\odot}$ and radius $R=10.29^{+1.01}_{-0.86}\leavevmode\nobreak\ \mathrm{km}$ (Mauviard et al., 2025b), and PSR J0437–4715 ($M=1.418\pm 0.037\,M_{\odot}$, $R=11.36^{+0.95}_{-0.63}\,\mathrm{km}$) (Choudhury et al., 2024a).
@@
Furthermore, tidal deformability, which describes how a NS deforms under the gravitational influence of its companion in a binary system, is also taken into account. This parameter is highly sensitive to the EOS and the compactness of the star (Hinderer, 2008; Hinderer et al., 2010). We therefore examine whether our hybrid model aligns with the tidal deformability constraint for a $1.4M_{\odot}$ NS, given as $\Lambda_{1.4}=190^{+390}_{-120}$ (Abbott et al., 2018).
@@
To evaluate the reliability of our model and explore how the DM parameters align with observational constraints on NS properties, both the pure hadronic and hybrid EOSs are used as inputs to the Tolman–Oppenheimer–Volkoff (TOV) equations (Tolman, 1939; Oppenheimer & Volkoff, 1939). This allows us to obtain the M–R relations corresponding to each EOS. The results are displayed in Fig. 4, where the observational M–R constraints are shown as color-coded regions for comparison.
In Fig. 4, the M–R curve for the pure hadronic EOS, which includes hyperons (DD2Y-T), is shown with a blue dashed line. The red dash-dotted line represents the hybrid EOS without any DM (S) particles (RIC_DD2Y-T). The solid colored lines show the full hybrid model that includes hyperons, S particles, and QM, where each color corresponds to a different S mass as indicated in the legend.
From the M–R curves, we can see that increasing the mass of S particles shifts the phase transition to QM to higher densities, meaning it happens at higher NS masses. This can be observed from the point where the curves start to deviate from the pure hadronic one. Because of the production of DM particles, the radius of the star changes compared to the model without S, bringing the results in better agreement with all the observational constraints. While all S masses satisfy the high-mass limit, the region corresponding to HESS J1731-347 is only consistent with the lowest S masses ($m_{\text{S}}\lesssim 1900$ MeV), due to the earlier deconfinement transition. The latest NICER M-R results for PSR J0614-3329, presented with maroon regions, indicate a measured mass of $\sim 1.4\,M_{\odot}$ having a radius comparable to HESS J1731-347. As also shown in the zoomed-in panel (a), all hybrid EOSs constructed using the RIC method and including S DM are found to intersect the $95\%$ credible region of PSR J0614-3329. In contrast, both the RIC_DD2Y-T and DD2Y-T models, which either exclude DM or represent purely hadronic matter, fail to meet this observational constraint. This highlights that the consistency of a model with such a compact object requires additional softening beyond what is provided by hyperons or QM alone. Therefore, the presence of an additional degree of freedom, such as S DM, can supply the necessary softening and make our hybrid model consistent with the most recent NICER results.
@@
We have also used the NICER measurements for PSR J0030+0451 from Vinciguerra et al. (2024), which represent a recent update to the analysis by Riley et al. (2019). Examining the cyan regions, the impact of the S becomes evident. In contrast to the RIC_DD2Y-T and DD2Y-T models, the EOSs that include S DM successfully pass through the $68\%$ credible region for PSR J0030+0451, indicating strong agreement with the most likely observational results.
@@
Figures 5 and 6 show how the presence of S particles affects the tidal deformability of NSs. In Fig. 5, we show how the dimensionless tidal deformability, $\Lambda$, changes with the mass of the star. The green vertical line shows the observational limit for a $1.4M_{\odot}$ NS from the GW170817 event (Abbott et al., 2018). Different lines in the plot correspond to different S particle masses, as shown in the legend.
@@
For completeness, Fig. 6 shows the allowed region in the $\Lambda_{1}$$\Lambda_{2}$ plane based on the GW170817 observation (Abbott et al., 2017). This figure also supports the upper limit of around 2040 MeV for the S particle mass. It can be seen that for $m_{\text{S}}\lesssim 2040$ MeV, the curves fall within the $90\%$ confidence region. In contrast, for heavier S particles, as well as for the pure hadronic and hybrid models without S, the curves fall outside the allowed region.
In conclusion, based on all the available multi-messenger data from NSs, we place a limit on the maximum allowed mass of the S particle. We find that the S mass should lie between 1850 and 2040 MeV to satisfy both nuclear and astrophysical constraints. Within this range, models that include hyperons, DM, and deconfined QM can all exist in the NS core. Among them, lighter S masses around 1900 MeV are especially favored, considering the M-R data from HESS J1731–347 and PSR J0437–4715.
Figure 5: The dimensionless tidal deformability, denoted as $\Lambda$, is presented as a function of stellar mass for the ensemble of EOS illustrated in Fig. 4. The green vertical line in the figure signifies the $\Lambda_{1.4}$ constraint derived from the low-spin prior analysis of GW170817, as reported in (Abbott et al., 2018).
Figure 6: Tidal deformability parameters $\Lambda_{1}$ and $\Lambda_{2}$ of the high- and low-mass components of the binary merger. The green region is the placed constraint on the tidal effects by the LIGO and Virgo collaboration from the GW170817
event (Abbott et al., 2018).

### v3

@@ 3 Equation of state characterization across phases and thermodynamic behavior
Figure 1: Pressure as a function of baryonic chemical potential for both hadronic phase and quark phase. The solid lines correspond to the hadronic phase with S for which the coupling strength of S remains constant ($x_{\text{S}}=0.03$), while different values of the mass of S are supposed. The QM EOSs shown with dotted lines are characterized by two key parameters: the diquark coupling ($\eta_{D}$) and the vector meson coupling ($\eta_{V}$), respectively. The hadronic EOS without S is also shown with the dashed blue line as DD2Y-T for comparison. The colored dots show the critical chemical potential in Eq. (8) at which the RIC has been applied.
@@
This study aims to explore the upper limit for the mass of S as a candidate for DM within the framework of the DD2Y-T model. To determine the allowed range for the mass of S, it is important to explore the behavior of matter at high densities and make sure the results agree with the observational constraints from NSs. It is essential to note that a mass shift of S particles, $x_{\text{S}}$, serves the purpose of mimicking the interaction with the surrounding medium. Given the postulation that S is a candidate for DM, it is reasonable to anticipate a predominantly weak interaction with the medium. Moore and Slatyer (2024) emphasized that for S particles to be a major DM candidate, their interactions must be extremely suppressed. Consequently, even though increasing the mass of S would inherently enable us to increase the value of $x_{\text{S}}$ based on our previous study (Shahrbaf et al., 2022), it is a deliberate choice to maintain $x_{\text{S}}$ as the coupling strength of the DM interaction with BM at its minimal value, specifically $x_{\text{S}}=0.03$ throughout the subsequent analyses. Although the chosen value of $x_{\text{S}}$ is small, Eq. (3) shows that increasing the baryon density still leads to a significant $\Delta m_{\text{S}}$, effectively mimicking the interactions between DM and BM in the dense cores of NSs. This simplifying approach allows us to systematically vary the mass of the DM candidate and examine its impact on the consistency of the hybrid star EOS, which includes hyperons in the hadronic phase, with current observational constraints.
@@
Therefore, we applied RIC, as outlined in Section 2.3, to model the phase transition from hadronic matter to deconfined QM. Within this approach, we studied the influence of the mass of S particles on the hybrid EOS and aim to determine the maximum allowed value for the mass of S.
@@
Figure 1 illustrates the pressure as a function of the baryonic chemical potential, showcasing both the hadronic phase and the quark phase. In the context of the hadronic phase, the coupling strength of S remains constant, as previously discussed, while different values of the mass of S with an increasing trend are supposed. One should notice that in comparison to the hadronic EOS without S (as represented by DD2Y-T, which includes nucleons and hyperons), the inclusion of S as a DM candidate results in an obvious softening of the EOS, although a repulsive interaction between DM and BM is considered. In the quark phase, we depict two various EOSs with specific parameters determined through BA, as outlined in reference Shahrbaf et al. (2023). Among the posterior-favored nlNJL parametrizations, we specifically select those that yield the earliest possible phase transition when combined with each hadronic EOS. This conservative choice ensures that deconfinement sets in as early as allowed by the model. Without this criterion, each hadronic EOS could, in principle, intersect with several different QM curves, leading to multiple possible transition points.
@@
The color dots show the crossing point between hadronic and quark lines, where the RIC with a negative value of $\Delta_{P}$ in Eq. (9) has been applied to correct the mischaracterization of a transition from hadronic matter to deconfined QM. As a result, at lower densities, the hadronic phase dominates, while at high densities, the quark phase remains a valid description. One should also point out that the RIC construction with $\Delta_{P}<0$ generates a peak in the squared speed of sound (as shown in Fig. 11 of Shahrbaf et al. (2022)), consistent with the phenomenologically established behavior observed in QCD-based studies.
@@
It is important to emphasize again that, for each hadronic EOS, we have selected the first intersection point with the QM EOS as the basis for implementing the phase transition construction. Although the two QM EOSs used in this study differ only slightly in stiffness, the softest three hadronic EOSs, corresponding to $m_{\text{S}}=$ 1890, 1895, and 1900 MeV, do not intersect with the stiffer of the two quark EOSs. Therefore, in the next section, where we perform a detailed scan over the mass of the S particle, these three soft hadronic models are combined with the first QM EOS (characterized by $\eta_{D}=0.70$, $\eta_{V}=0.10$). The remaining hadronic models are matched with the second, slightly stiffer QM EOS (with $\eta_{D}=0.70$, $\eta_{V}=0.11$). For both parametrizations, the squared speed of sound is inferred to be $c_{s}^{2}=0.44$.
@@
[…]
Figure 3: Pressure as a function of energy density for the baseline hadronic EOSs (DD2 and DD2Y-T) and the constructed hybrid EOSs. The yellow shaded region and dashed green line indicate the constraints from Hebeler et al. (2013) and Miller and others (2021), respectively.
Figure 3, shows the pressure as a function of energy density for the baseline hadronic EOSs (DD2 and DD2Y-T) and the constructed hybrid EOSs. All employed and constructed EOSs are consistent with the Hebeler et al. (Hebeler et al., 2013) constraint (yellow region) and largely fall within the narrower Miller et al. (Miller and others, 2021) region (between dashed green lines), with only slight deviations at higher S-particle masses. At low energy densities, where hadronic matter dominates, the hybrid EOSs with DM are softer than the reference hadronic lines (DD2: black dash-dotted, DD2Y-T: light red lines) and also the hybrid RIC-DD2Y-T (violet) line, with stiffness increasing for higher S masses. At high energy densities, where QM governs, all hybrid EOSs converge, as both quark EOS sets share the same stiffness and speed of sound, rendering the high-density behavior almost insensitive to the S DM mass. At this region, the hybrid EOSs with DM (RIC-DD2Y-T+S) remain softer than DD2 but stiffer than DD2Y-T and RIC-DD2Y-T.
@@
Figure 4: Fraction of S particles as a DM candidate in the hadronic phase calculated using the DD2Y-T model for various S particle masses. The fractions are shown from the respective onset densities up to densities beyond the central density of NSs, illustrating how the S fraction evolves within the purely hadronic phase. The black and red dash-dotted lines indicate the onset densities for the transition to deconfined QM, corresponding to $\mu_{H}$ for $m_{S}=1890$ MeV and $m_{S}=2054$ MeV, respectively.
@@
Regarding the production of S particles inside the hybrid star modeled by the approach described above, Fig. 4 shows the variation of the DM particle fraction as a function of baryon number density for different S-particle masses. In general, lighter S particles result in a higher DM fraction at a given density. Increasing the mass of S shifts the DM onset to higher densities, leading to a smaller fraction of heavier S particles being produced in stellar matter. For a fixed S mass, the DM fraction increases as the system becomes denser. However, at high densities, around $n_{B}\approx 1fm^{-3}$, the DM fraction becomes roughly the same across all considered S masses shown in the legend. The vertical black and dash-dotted red lines indicate the deconfinement onset densities, corresponding to $\mu_{H}$ for $m_{S}=1890$ MeV and $m_{S}=2054$ MeV, respectively, marking the chemical potentials at which some degrees of phase transition to QM occur. It is worth noting that the microscopic evolution of particle fractions within the mixed phase is not explicitly determined in our approach. The onset of the transition is specified by $\mu_{H}$, and the maximum DM fraction is reached at this threshold density for each stellar mass. Consequently, the maximum DM fraction allowed in our models varies from about $12\%$ for the highest S-particle mass to $15\%$ for the lightest one. Although the DM fractions vary along the curves, from zero up to about $30\%$ at very high densities within the pure hadronic phase, the maximum S fraction inside hybrid stars changes only within a narrow range of $12$-$15\%$. This behavior indicates that increasing the DM particle mass not only shifts the DM onset to higher densities but also delays the deconfinement transition, pushing it to even higher densities, as also seen in Fig. 1.
4 Neutron star constraints on the mass of dark matter particles
@@
Figure 5: Mass-radius relations for NSs with a purely hadronic core (dashed blue line), hybrid stars without DM (dash-dotted red line), and hybrid stars including DM for different S particle masses (solid colored lines). The colored regions represent observational constraints for comparison. The 1-$\sigma$ M-R constraint from the NICER analysis of PSR J0740+6620 is shown in turquoise (Dittmann et al., 2024). The updated $95\%$ and $68\%$ credible NICER results for PSR J0030+0451 are displayed as nested light and dark cyan regions (Vinciguerra et al., 2023). The hatched magenta and green regions show the high-mass BW pulsar PSR J0952-0607 (Romani et al., 2022) and PSR J0348+0432 (Antoniadis and others, 2013), respectively. The credible intervals for HESS J1731-347 are shown as nested green regions, representing the $68\%$ (inner) and $90\%$ (outer) levels based on X-ray spectral modeling (Doroshenko et al., 2023). The M-R constraints for PSR J0437-4715 are represented by overlapping dark and light blue regions, corresponding to the $68\%$ (inner) and $90\%$ (outer) credible intervals (Choudhury et al., 2024). The nested light and dark maroon regions show the most recent NICER analysis for the mass and radius of PSR J0614-3329, $95\%$ and $68\%$ credible results, respectively (Mauviard et al., 2025). The zoomed-in region (a) highlights that all RIC EOSs including S DM pass through the $95\%$ credible region of PSR J0614–3329, whereas the RIC_DD2Y-T and DD2Y-T curves do not satisfy this constraint. The zoomed-in region (b) illustrates that for S masses larger than $1930$ MeV, the NS radius lies outside the $68\%$ credible region for PSR J0437-4715.
@@
As explained in the previous sections, recent multimessenger observations of NSs offer a valuable chance to explore the structure of high-density matter, which may contain exotic components such as hyperons, QM, or even DM. In this section, based on the astrophysical constraints of NSs, we investigate our hybrid EOS and determine the allowed maximum mass for the S particle. For this purpose, the latest M-R measurements by NICER telescope have been applied, which provide reliable constraints for the radius of NSs with masses around $2M_{\odot}$ and $1.4M_{\odot}$ (Miller and others, 2019; Riley and others, 2019; Miller and others, 2021; Riley and others, 2021; Choudhury et al., 2024; Mauviard and others, 2025). Since all of our hybrid EOSs pass well through the center of the NICER constraint region for PSR J0740+6620 with a mass of $2M_{\odot}$, we now focus primarily on the most recent results for the pulsars PSR J0614-3329 which has a measured mass of $M=1.44^{+0.06}_{-0.07}\,M_{\odot}$ and radius $R=10.29^{+1.01}_{-0.86}\penalty 10000\ \mathrm{km}$ (Mauviard and others, 2025), and PSR J0437-4715 ($M=1.418\pm 0.037\,M_{\odot}$, $R=11.36^{+0.95}_{-0.63}\,\mathrm{km}$) (Choudhury et al., 2024).
@@
Furthermore, tidal deformability, which describes how a NS deforms under the gravitational influence of its companion in a binary system, is also taken into account. This parameter is highly sensitive to the EOS and the compactness of the star (Hinderer, 2008; Hinderer et al., 2010). We therefore examined whether our hybrid model aligns with the tidal deformability constraint for a $1.4M_{\odot}$ NS, given as $\Lambda_{1.4}=190^{+390}_{-120}$ (Abbott and others, 2018).
@@
To evaluate the reliability of our model and explore how the DM parameters align with observational constraints on NS properties, both the pure hadronic and hybrid EOSs are used as inputs to the Tolman–Oppenheimer–Volkoff equations (Tolman, 1939; Oppenheimer and Volkoff, 1939). This allows us to obtain the M–R relations corresponding to each EOS. The results are displayed in Fig. 5, where the observational M–R constraints are shown as color-coded regions for comparison.
In Fig. 5, the M–R curve for the pure hadronic EOS, which includes hyperons (DD2Y-T), is shown with a dashed blue line. The dash-dotted red line represents the hybrid EOS without any DM (S) particles (RIC_DD2Y-T). The solid colored lines show the full hybrid model that includes hyperons, S particles, and QM, where each color corresponds to a different S mass as indicated in the legend.
From the M–R curves, we can see that increasing the mass of S particles shifts the phase transition to QM to higher densities, meaning it happens at higher NS masses. This can be observed from the point where the curves start to deviate from the pure hadronic one. Because of the production of DM particles and the resulting softening, the radius of the star changes compared to the model without S, bringing the results in better agreement with all the observational constraints. The softening induced by DM acts mainly in the hadronic phase, which reduces radii at $\sim 1.4M_{\odot}$. By contrast, the maximum mass is governed predominantly by the high-density quark EOS, which is essentially unchanged across our cases because both quark-matter parametrizations result in the same $c_{s}^{2}$. Consequently, the hybrid EOSs with DM yield comparable $M_{\max}$. The onset density of the phase transition also matters. For example, the RIC-DD2Y-T EOS (without DM) has a late transition, retains a larger hadronic fraction at high density, and therefore attains a smaller $M_{\max}$. While all S masses satisfy the high-mass limit, the region corresponding to HESS J1731-347 is only consistent with the lowest S masses ($m_{\text{S}}\lesssim 1900$ MeV), due to the earlier deconfinement transition. The latest NICER M-R results for PSR J0614-3329, presented with maroon regions, indicate a measured mass of $\sim 1.4\,M_{\odot}$ having a radius comparable to HESS J1731-347. As also shown in the zoomed-in panel (a), all hybrid EOSs constructed using the RIC method and including S DM are found to intersect the $95\%$ credible region of PSR J0614-3329. In contrast, both the RIC_DD2Y-T and DD2Y-T models, which either exclude DM or represent purely hadronic matter, fail to meet this observational constraint. This highlights that the consistency of a model with such a compact object requires additional softening beyond what is provided by hyperons or QM alone. Therefore, the presence of an additional degree of freedom, such as S DM, can supply the necessary softening and make our hybrid model consistent with the most recent NICER results.
@@
We have also used the NICER measurements for PSR J0030+0451 from Vinciguerra and others (2024), which represent a recent update to the analysis by Riley and others (2019). Examining the cyan regions, the impact of the S becomes evident. In contrast to the RIC_DD2Y-T and DD2Y-T models, the EOSs that include S DM successfully pass through the $68\%$ credible region for PSR J0030+0451, indicating strong agreement with the most likely observational results. We note that our hybrid EOSs remain consistent with the NICER measurement of PSR J1231-1411 (Salmi and others, 2024a), although this object is still under debate due to its complex pulse profile; for this reason, we did not include it in our M-R diagram.
@@
Figures 6 and 7 show how the presence of S particles affects the tidal deformability of NSs. In Fig. 6, we show how the dimensionless tidal deformability, $\Lambda$, changes with the mass of the star. The vertical green line shows the observational limit for a $1.4M_{\odot}$ NS from the GW170817 event (Abbott and others, 2018). Different lines in the plot correspond to different S particle masses, as shown in the legend.
@@
For completeness, Fig. 7 shows the allowed region in the $\Lambda_{1}$$\Lambda_{2}$ plane based on the GW170817 observation (Abbott and others, 2017). This figure also supports the upper limit of around 2040 MeV for the S particle mass. It can be seen that for $m_{\text{S}}\lesssim 2040$ MeV, the curves fall within the $90\%$ confidence region. In contrast, for heavier S particles, as well as for the pure hadronic and hybrid models without S, the curves fall outside the allowed region.
In conclusion, based on all the available multi-messenger data from NSs, we place a limit on the maximum allowed mass of the S particle. We find that the S mass should lie between 1885 and 2040 MeV to satisfy both nuclear and astrophysical constraints. Within this range, models that include hyperons, DM, and deconfined QM can all exist in the NS core. Among them, lighter S masses around 1900 MeV are especially favored, considering the M-R data from HESS J1731–347 and PSR J0437-4715.
Figure 6: Dimensionless tidal deformability, denoted as $\Lambda$, presented as a function of stellar mass for the ensemble of EOSs illustrated in Fig. 5. The vertical green line in the figure signifies the $\Lambda_{1.4}$ constraint derived from the low-spin prior analysis of GW170817, as reported in (Abbott and others, 2018).
Figure 7: Tidal deformability parameters $\Lambda_{1}$ and $\Lambda_{2}$ of the high- and low-mass components of the binary merger. The green region is the placed constraint on the tidal effects by the LIGO and Virgo collaboration from the GW170817 event. (Abbott and others, 2018).